# Functional-level Uncertainty Quantification for Calibrated Fine-tuning on LLMs

### v3

@@ 5.1 In-distribution Performance
As shown in Figure 3, UQ4CT achieves notable gains in uncertainty calibration across diverse tasks, while maintaining competitive accuracy (ACC) relative to baseline approaches. For instance, on BoolQ and ClimateQA tasks, UQ4CT attains accuracy rates of $73.5\%$ and $63.3\%$, demonstrating that improved uncertainty quantification does not come at the expense of predictive performance.
The most significant improvements are seen in reduced Expected Calibration Error (ECE). UQ4CT consistently lowers ECE by over $25\%$ on average across benchmarks, and unlike other approaches, it continues to perform well even on challenging datasets such as ARC-C, achieving an ECE of $6.1$.
To further validate our method, we include results from fine-tuning both LLaMA-2-7B (main text) and Mistral-7B (Appendix A.3). Across both models, UQ4CT delivers substantial and consistent improvements in uncertainty calibration. These results underscore the practical value of UQ4CT, particularly in scenarios where reliable uncertainty estimates are crucial, such as safety-critical applications. A key advantage of UQ4CT is that it incorporates uncertainty calibration directly during fine-tuning, incurring minimal computational overhead compared to other uncertainty quantification (UQ) methods, which often require costly repetitive sampling or post-hoc adjustments.
@@ 5.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, a). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios.
@@ 5.3 Ablation Studies
$\beta$ 00 $0.2$ $0.5$ $0.8$ $1$ $1.2$ $1.5$ $1.8$ $2$
ACC $\uparrow$ $75.5_{2.91}$ $76.0_{0.6}$ $75.9_{0.31}$ $76.6_{0.4}$ $77.3_{1.36}$ $\mathbf{77.9_{0.6}}$ $77.2_{0.91}$ $76.7_{0.82}$ $76.6_{0.92}$
ECE $\downarrow$ $12.9_{1.99}$ $7.47_{0.78}$ $7.82_{0.93}$ $5.94_{1.16}$ $\mathbf{5.0_{1.15}}$ $6.49_{2.65}$ $9.19_{0.86}$ $8.11_{1.43}$ $6.65_{1.85}$
Sensitivity Test on Calibration Term. To further understand the effectiveness of the calibration loss, we perform a sensitivity test of the coefficient $\beta$ in Equation 14. This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities. We evaluate $\beta$ values ranging from 00 to $2$, where $\beta=0$ resembles the original MixLoRA method.
@@ A.1 Theoretical Derivation of the Method
¯$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
@@
¯$\bar{\mathcal{L}}_{\mathrm{cal}}$ $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}$
@@ A.3 Experimental Results with Mistral-7B
ACC $\uparrow$ UQ4CT $73.5_{0.52}$ $76.6_{1.30}$ $52.8_{1.77}$ $77.3_{1.36}$ $63.3_{1.74}$
UQ4CT_Incremental $72.0_{0.19}$ $75.4_{0.81}$ $54.6_{0.95}$ $77.6_{0.43}$ $60.2_{3.17}$
ECE $\downarrow$ UQ4CT $\mathbf{2.3_{0.82}}$ $\mathbf{6.0_{0.2}}$ $\mathbf{6.1_{1.11}}$ $\mathbf{5.0_{1.15}}$ $\mathbf{8.1_{0.52}}$
UQ4CT_Incremental $4.0_{0.16}$ $9.8_{1.51}$ $13.8_{2.08}$ $10.3_{1.73}$ $12.2_{0.88}$
Table 4: Performance comparison of different methods fine-tuned with Mistral-7B across $4$ common sense reasoning tasks and a domain-specific task. UQ4CT shows significant ECE improvements while maintaining high accuracy.
@@ A.4 Incremental Weighting on Calibration Term
Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data. Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration. We conduct this study by incrementally increase $\beta$ from 00 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
@@ A.6 Training Details
We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability. For each expert, we use $rank=16$ and $alpha=32$. We use batch size of $16$ to train our model. For climate task, we set the learning rate to $45e-$ 4 and dropout rate to $0.1$ to incorporate the small dataset size. For other tasks, we use $42e-$ 4 as our learning rate with dropout $0.05$. We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training. Our model is trained on A100 GPU, with 20GB GPU memory consumption per task. Training time is from $25$ to $50$ minutes depending on the task.

### v4

@@ 5.1 In-distribution Performance
As shown in Table 1, UQ4CT achieves notable gains in uncertainty calibration across diverse tasks, while maintaining competitive accuracy (ACC) relative to baseline approaches. For instance, on OBQA and ClimateQA tasks, UQ4CT attains accuracy rates of $88.4\%$ and $79.9\%$, demonstrating that improved uncertainty quantification does not come at the expense of predictive performance.
The most significant improvements are seen in reduced Expected Calibration Error (ECE). UQ4CT consistently lowers ECE by over $25\%$ on average across benchmarks, and unlike other approaches, it continues to perform well even on challenging datasets such as ARC-C, achieving an ECE of $4.4$.
To further validate our method, we include results from fine-tuning both LLaMA-3.1-8B (main text) and Mistral-7B (Appendix A.3). Across both models, UQ4CT delivers substantial and consistent improvements in uncertainty calibration. These results underscore the practical value of UQ4CT, particularly in scenarios where reliable uncertainty estimates are crucial, such as safety-critical applications. A key advantage of UQ4CT is that it incorporates uncertainty calibration directly during fine-tuning, incurring minimal computational overhead compared to other uncertainty quantification (UQ) methods, which often require costly repetitive sampling or post-hoc adjustments.
@@ 5.2 Performance Under Distribution Shift
Due to the sparse nature of the fine-tuning data, real world deployment of LLMs often requires the model to be robust to out-of-distribution knowledge (Ouyang et al., 2022; Touvron et al., 2023d, c). Therefore, we evaluate the performance of UQ4CT along with other baseline models fine-tuned on the OBQA dataset under smaller and larger distribution shift scenarios.
@@ 5.3 Ablation Studies
$\beta$ 0 $0.2$ $0.5$ $0.8$ 11 $1.2$ $1.5$ $1.8$ 22
[…]
ECE $\downarrow$ $12.7_{1.92}$ $7.35_{0.75}$ $7.69_{0.89}$ $5.82_{1.12}$ $\mathbf{3.34_{1.60}}$ $6.31_{2.58}$ $9.03_{0.82}$ $7.96_{1.38}$ $6.52_{1.79}$
Sensitivity Test on Calibration Term. To further understand the effectiveness of the calibration loss, we perform a sensitivity test of $\beta$ in Equation 14. This evaluates how our proposed calibration of parameter mixtures affect the overall model prediction and uncertainty quantification capabilities. We evaluate $\beta$ values ranging from 0 to $2$, where $\beta=0$ resembles the original MixLoRA method.
@@ A.1 Theoretical Derivation of the Method
$\bar{\mathcal{L}}_{\mathrm{cal}}=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}.$
@@
$\bar{\mathcal{L}}_{\mathrm{cal}}$ $=\mathbb{E}_{(x,y^{*})\sim\mathcal{D}}\left(\mathbbm{1}\{\mathrm{MixLoRA}(x)=y^{*}\}-\mathrm{FLU}(x)\right)^{2}$
@@ A.3 Experimental Results with Mistral-7B
ACC $\uparrow$ UQ4CT $9.17_{1.33} 88$$8.66_{0.20}$ $79.60_{1.21}$ $88.40_{0.35}$ $79.97_{0.85}$
UQ4CT_Incremental $87.33_{0.24}$ $87.15_{0.95}$ $0.84_{1.03} 88$$8.53_{0.48}$ $75.87_{2.89}$
ECE $\downarrow$ UQ4CT $\mathbf{1.79_{0.43}}$ $\mathbf{3.97_{0.78}}$ $\mathbf{4.43_{0.82}}$ $\mathbf{3.34_{1.60}}$ $\mathbf{9.36_{2.77}}$
UQ4CT_Incremental $3.10_{0.18}$ $6.55_{1.42}$ $10.02_{1.95}$ $6.87_{1.68}$ $14.16_{0.91}$
Table 5: Performance comparison of different methods fine-tuned with Mistral-7B across 44 common sense reasoning tasks and a domain-specific task. UQ4CT shows significant ECE improvements while maintaining high accuracy.
@@ A.4 Incremental Weighting on Calibration Term
Due to the random initialization of LoRA experts, the predictions during early fine-tuning stage are likely to be incorrect as the model has little knowledge on the functional relationships regarding the data. Thus, it is intuitive to incrementally increase the weight parameter $\beta$ over the calibration term $\mathcal{L}_{\mathrm{cal}}$ in the training loss for the LoRA experts to learn before calibration. We conduct this study by incrementally increase $\beta$ from 0 to $1$ within $50$ gradient steps during the early stage of fine-tuning:
@@ A.6 Training Details
We train our model with total of $8$ LoRA experts, and select $2$ experts with the highest probability. For each expert, we use $rank=16$ and $alpha=32$. We use batch size of $16$ to train our model. For climate task, we set the learning rate to 5e45e-4 and dropout rate to $0.1$ to incorporate the small dataset size. For other tasks, we use 2e42e-4 as our learning rate with dropout $0.05$. We use AdamW as our optimizer and a cutoff length of $512$ for prompts during training. Our model is trained on A100 GPU, with 20GB GPU memory consumption per task. Training time is from $25$ to $50$ minutes depending on the task.